OpenStudy (jacobbirddog):

Please help... 5 questions (i did 23 of them) I cant figure them out!!! Fan and medal!

OpenStudy (jacobbirddog):

Simplify: (3a – 4b)(a + b) A. 3a2– ab – 4b2 B. 3a2 – 7ab – 4b2 C. 3a2 – 12ab – 4b2 D. 3a2 + ab – 4b2

OpenStudy (anonymous):

3a*a=3a^2 -4b*a=-4ba 3a*b=3ab -4b*b=-4b^2 We combine like terms, and the final answer is 3a^2-ab-4b^2

OpenStudy (jacobbirddog):

i kept on getting b -_- thank you!

OpenStudy (anonymous):

You're welcome, do you understand how I got my answer?

OpenStudy (jacobbirddog):

well i was doing it wrong obviously but i do not quite understand please explain further.

OpenStudy (anonymous):

Okay, so if we had a problem with multiplying (a+b)(c-d) The first step would be a*c=ac Then we would multiply b*c=bc Then we would multiply a*-d=-ad Then we would multiply b*-d=-bd Then we combine like terms and put them all together, but there are no like terms in this example so we end up with ac+bc-ad-bd

OpenStudy (jacobbirddog):

ahhh i see thankyou for explainting it in detail :)

OpenStudy (jacobbirddog):

by the way im japanese and im learning english... am i doing ok?

OpenStudy (anonymous):

Yes, I had no idea English was not your first language!

OpenStudy (jacobbirddog):

thank you!

OpenStudy (jacobbirddog):

i am translating these problems by the way

OpenStudy (jacobbirddog):

Simplify. (4a – 3)2 A. 8a2 – 24a + 6 B. 162a – 12 + 9 C. 16a2 – 9 D. 16a2 – 24a + 9

OpenStudy (jacobbirddog):

Simplify. (4a – 3)^2 A. 8a^2 – 24a + 6 B. 16^2a – 12 + 9 C. 16a^2 – 9 D. 16a^2 – 24a + 9

OpenStudy (jacobbirddog):

i added carats so its easier

OpenStudy (anonymous):

Okay, well the exponent goes to each value inside the parentheses so it will look like this: 4^2x^2-3^2 We then simplify the equation 4^2=16 -3^2=9 (because we are multiplying -3 by -3 it becomes positive 9 because a negative * negative=positive Our simplified equation will be 16a^2+9

OpenStudy (jacobbirddog):

i got c so i think i got it right i mean, i worked the problem 2 times and got c

OpenStudy (anonymous):

Yeah you're right it is 16a^2-9, sorry.

OpenStudy (jacobbirddog):

thanks!

OpenStudy (jacobbirddog):

Simplify. (x – 5)(x2 – 2x – 6) A. x3 + x2 – 4x + 30 B. x3 – 5x2 + 10x + 30 C. x3 – 7x2 + 16x + 30 D. x3 – 7x2 + 4x + 30

OpenStudy (anonymous):

Going through the same process that I showed you, I got x^3-7x^2+4x+30

OpenStudy (jacobbirddog):

i got a so i was checking

OpenStudy (anonymous):

I am pretty sure that the answer is D.

OpenStudy (jacobbirddog):

mk

OpenStudy (jacobbirddog):

one last q

OpenStudy (jacobbirddog):

Find the volume of the box. Use the formula V = lwh. http://static.k12.com/calms_media/media/330000_330500/330323/1/576833d32994807ff1480a750d977fb89e6b2e71/MS_ALG_S2_03_06_test_Q5_question.gif A. 6x3 – 2 B. 6x3 + 11x2 – 3x – 2 C. 6x2– x – 1 D. 6x2 – 13x2 – 3x – 2

OpenStudy (anonymous):

I got 6x^3+11x^2-3x-2 (letter B.)